what is the H3O+ concentration of a solution with a pH of 2
1
Expert's answer
2017-04-26T08:08:14-0400
what is the H3O+ concentration of a solution with a pH of 2
Solution:
[H3O+] = [H+] = 10-pH;
[H3O+] = 10-2 = 0.01 M.
Answer: 0.01 M.